OpenStudy (dugalde):

convert z=(-5 sqrt 3)/2+(5/2)i and w=1+(sqrt3)i to polar form

OpenStudy (anonymous):

HI!!

OpenStudy (anonymous):

learned that from @misty1212

OpenStudy (anonymous):

you need two numbers to write \[-\frac{5\sqrt3}{2}+\frac{5}{2}i\] in polar form \[r\left(\cos(\theta)+i\sin(\theta)\right)\] namely \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r\) is the easiest it is \[r=\sqrt{a^2+b^2}\] which in your case is \[r=\sqrt{(\frac{5\sqrt3}{2})^2+(\frac{5}{2})^2}\] let me know what you get

OpenStudy (dugalde):

I get 5

OpenStudy (anonymous):

yeah me too

OpenStudy (anonymous):

now you need \(\theta\) \[\cos(\theta)=\frac{a}{r}\] which in your case is conveniently \[-\frac{\sqrt3}{2}\]

OpenStudy (dugalde):

i got -8.66

OpenStudy (anonymous):

also \[\sin(\theta)=\frac{b}{r}=\frac{1}{2}\] find an angle \(\theta\) with \[\cos(\theta)=-\frac{\sqrt3}{2}\] and \[\sin(\theta)=\frac{1}{2}\]

OpenStudy (anonymous):

use radians, not decimals

OpenStudy (dugalde):

don't know how to do that

OpenStudy (anonymous):

it is a really really familiar point on the unit circle \[\left(-\frac{\sqrt3}{2},\frac{1}{2}\right)\]

OpenStudy (anonymous):

you got a unit circle cheat sheet?

OpenStudy (anonymous):

now you do

OpenStudy (dugalde):

5pi/6

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

polar form \[5\left(\cos(\frac{5\pi}{6})+i\sin(\frac{5\pi}{6})\right)\]

OpenStudy (dugalde):

what about w=1+(sqrt 3)i

OpenStudy (anonymous):

repeat the process above \[r=\sqrt{a^2+b^2}\] and \[\cos(\theta)=\frac{a}{r}\\ \sin(\theta)=\frac{b}{r}\] finds you \(r\) and \(\theta\)

OpenStudy (dugalde):

r=2

OpenStudy (dugalde):

theta=60 degrees

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (dugalde):

so what would the polar form be

OpenStudy (xapproachesinfinity):

use radians not degrees

OpenStudy (xapproachesinfinity):

what is 60 in radians

OpenStudy (dugalde):

1.04

OpenStudy (xapproachesinfinity):

polar \[w=2(\cos \theta+i\sin \theta) \] \[\theta =60 ~~degree \]

OpenStudy (xapproachesinfinity):

no no decimal my friend see what satellite did with the first use radians

OpenStudy (dugalde):

don't know who to do

OpenStudy (xapproachesinfinity):

see the cheat sheet

OpenStudy (dugalde):

pi/3

OpenStudy (xapproachesinfinity):

yes